Assume that the mean hourly cost to operate a commercial airplane follows the normal distribution with a mean of $1900 per hour and a standard deviation of $250.
What is the maximum operating cost for the lowest 5% of the airplanes? (Round z-score computation to 2 decimal places and the final answer to 1 decimal place.)
Maximum operating cost $

Answer by Boreal(15235)

z.05=-1.65
z=(x-mean)/sd=(x-1900)/250=-1.65
x-1900=-412.5
x=$1487.5